OpenStudy (safyousuff97):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (safyousuff97):

@MTALHAHASSAN2

OpenStudy (mtalhahassan2):

@safyousuff97 check out this link: http://new.openstudy.com/updates/5255dc2de4b037f0588761ef

OpenStudy (mtalhahassan2):

@baru

OpenStudy (baru):

i'll give a hint: there is a $65 base price, and a daily rent of $80

OpenStudy (safyousuff97):

so thats the domain and range

OpenStudy (baru):

no its not, and the question doesn't ask you for range and domain

OpenStudy (safyousuff97):

what is the first part asking then ?

OpenStudy (thebossofme191):

I'm here

OpenStudy (thebossofme191):

Lol like I'd be much help

OpenStudy (safyousuff97):

ok thanks alot :)

OpenStudy (safyousuff97):

will u guys be able too help?

OpenStudy (thebossofme191):

Hmmmmmm

OpenStudy (lexi_loves):

Im stumped.

OpenStudy (safyousuff97):

me too

OpenStudy (thebossofme191):

@tatumlee

OpenStudy (thebossofme191):

Alright I'm stumped on this too sry

OpenStudy (safyousuff97):

its ok at least u tried to help

OpenStudy (thebossofme191):

I'm sry I jus couldn't get it

OpenStudy (safyousuff97):

its ok

OpenStudy (thebossofme191):

OK bye I'm leave this now hope u get!it

OpenStudy (safyousuff97):

k hopefully

OpenStudy (safyousuff97):

@AaronAndyson

OpenStudy (maddie82502):

I believe there is a set price before renting. y = 65 + 80x I have no idea how I got it, I was just playing around with the numbers and it popped into my head. I know, weird. But, when we plug in, 5 = x. y = 65 + 80(5) y = 65 + 400 y = 465 When he rents for 5 days, he pays $465. Or when we plug in 7 = x y = 65 + 80(7) y = 65 + 560 y = 625 When he rents for 7 days, he paws $625. I hope that makes sense. ^^

OpenStudy (safyousuff97):

oh ok but which part do they all go in?

OpenStudy (safyousuff97):

maddie?

OpenStudy (maddie82502):

ok try dividing by seven

OpenStudy (maddie82502):

its part A bc each day costs $93 to rent so it would be y=x*93